OpenStudy (anonymous):

What is the slope of a line that is parallel to the line with the following equation: 4x + 6y = 8.

OpenStudy (anonymous):

-2/3

jimthompson5910 (jim_thompson5910):

4x + 6y = 8 6y = -4x+8 y = -4/6x+8/6 y=-2/3x+4/3 So the slope of this line is -2/3 Which means that the slope of any parallel line is also -2/3

OpenStudy (vijay):

-4/6

OpenStudy (saifoo.khan):

HEYY!

OpenStudy (anonymous):

vijay, so simplified it would be -2/3?

OpenStudy (vijay):

hmm

OpenStudy (anonymous):

Hey Saifoo!!!

jimthompson5910 (jim_thompson5910):

yes, -4/6 = -2/3

OpenStudy (saifoo.khan):

im sorry about yesterday! there was electricity failure @ my place!

OpenStudy (vijay):

oooho, which place

OpenStudy (anonymous):

Oh, that's no problem. I really enjoyed talking with you.

OpenStudy (saifoo.khan):

vijay, u knw Indo-pak problems... :(

OpenStudy (anonymous):

LOL!!

OpenStudy (saifoo.khan):

i enjoyed too! :)

OpenStudy (anonymous):

You educated me on some things I have been wondering about for a long time. I appreciate that.

OpenStudy (saifoo.khan):

Oh cool.. which things??

OpenStudy (anonymous):

Not just algebra things but your religion. It fascinates me.

OpenStudy (vijay):

indo ??

OpenStudy (saifoo.khan):

INDO-PAK = Indian & Pakistani Subcontinent.

OpenStudy (saifoo.khan):

Yeah.. Coool. :)

OpenStudy (anonymous):

You helped me to understand many things that I didn't know.

OpenStudy (vijay):

oh, south asia

OpenStudy (saifoo.khan):

Yea, are u an indian?

OpenStudy (saifoo.khan):

welll thts really nice! :)

OpenStudy (vijay):

nope,

OpenStudy (anonymous):

brb

OpenStudy (vijay):

i was born in London, my parents are Sri Lankans

OpenStudy (saifoo.khan):

Oh cool.. but stilll u r my neighbor.. lol.

OpenStudy (vijay):

hmm, but i dont like to say srilankan, becoz they r under the war crimes.

OpenStudy (saifoo.khan):

war crimes as in?

OpenStudy (anonymous):

sorry, I lost my internet connection!

OpenStudy (anonymous):

I feel so bad for you both.

OpenStudy (saifoo.khan):

why?o.O

OpenStudy (anonymous):

you were talking about being at war.

OpenStudy (anonymous):

I do not know what that is like and I feel bad.

OpenStudy (saifoo.khan):

oh hmmm..

OpenStudy (anonymous):

The last war this country had within itself was the civil war and that was back in the 1800's...long before I was in existence so I can't imagine what it must be like.

OpenStudy (anonymous):

Ohhh. thanks for the medals!!

OpenStudy (saifoo.khan):

LOl. in some parts od sri lanka there is civil war.. about which he is talking about.. np.

OpenStudy (anonymous):

For some reason I keep losing my internet connection.

OpenStudy (saifoo.khan):

and that reason is?

OpenStudy (anonymous):

I think it is because of the heat. It has been horribly hot for the past 3 weeks.

OpenStudy (saifoo.khan):

Ohh, might be.

OpenStudy (anonymous):

The heat index has been ranging from 101 to 113 degrees.

OpenStudy (anonymous):

It is unbearable.

OpenStudy (anonymous):

Thank God for A/C...LOL

OpenStudy (saifoo.khan):

yea, this same range here is in start of June.

OpenStudy (saifoo.khan):

Haha, hmmm..

OpenStudy (anonymous):

Oops, sorry. LOL

OpenStudy (anonymous):

It has not been this hot for so many days in a row since 1942.

OpenStudy (saifoo.khan):

Ohh,

OpenStudy (saifoo.khan):

soo?

OpenStudy (anonymous):

Well, I have to go fold some laundry so I will have to go now. I will talk to you soon. Sooo? what? LOL

OpenStudy (saifoo.khan):

Alright... Lol, i wanted to ask of there are any more question.

OpenStudy (saifoo.khan):

questions*